OpenStudy (rebecca1588):

What is the solution to the system of equations? x+y=5 x-2y=2 ( , )<<<<

OpenStudy (anonymous):

By substitution:\[x+y=5 <=> x=5-y\] Insert it in 2nd equation: \[(5-y)-2y=2 <=> y=1\] Insert y in first: \[x+1=5 <=> x=4\] You are correct.

OpenStudy (rebecca1588):

Thanks sooo much!!! :)